# TESSERACT Collaboration

### v2

@@ VI Acknowledgments
Supplementary Materials Section H Noise Comparison to Other Experiments
Here, we describe comparisons of the excess noise observed in the two detectors described in this work to other detectors.
Detector Name Dimensions Holding Correlated Noise Day of Run Noise / Mass Corrected Noise / Mass
(W2/Hz) (W2/(Hz g)) (W2/(Hz g))
This work, 4 mm 1 cm2 $\times$ 4 mm HangingAnthony-Petersen et al. (2024) $3.6\times 10^{-36}$ 12 $3.8\times 10^{-36}$ $3.8\times 10^{-36}$
This work, 1 mm 1 cm2 $\times$ 1 mm HangingAnthony-Petersen et al. (2024) $7.6\times 10^{-37}$ 12 $3.2\times 10^{-36}$ $3.2\times 10^{-36}$
SPICE 1%  [^1] This detector had degraded phonon collection efficiency and therefore lower observed noise due to accidental etching of its W TESs. We correct for this by scaling the noise up by the ratio of the phonon collection efficiencies measured for this work and for the SPICE 1% detector, i.e. (39.1%/28.9%) = 1.33. Anthony-Petersen et al. (2025) 1 cm2 $\times$ 1 mm HangingAnthony-Petersen et al. (2024) $1.0\times 10^{-36}$ 7 $4.4\times 10^{-36}$ $3.1\times 10^{-36}$
CPD v1222Single channel detector, noise reduced by a factor of two for comparison to two channel detectors. Fink et al. (2021); Alkhatib et al. (2021) 7.6 cm Ø$\times$ 1 mm Cirlex clamps $4.2\times 10^{-35}$ Unknown $4.0\times 10^{-36}$ $\sim 2.8\times 10^{-36}$
CPD v2333Not previously published, an updated version of the detector described in Ref. Fink et al. (2021). Single channel detector, noise reduced by a factor of two for comparison to two channel detectors. 7.6 cm Ø$\times$ 1 mm Sapphire clamps 4.7 ×$\times 10^{-35}$ 6 $4.5\times 10^{-36}$ $2.9\times 10^{-36}$
Table 2: Correlated noise amplitude comparison in different low threshold phonon detectors. The two single channel CPD detectors have their noise reduced by a factor of two to correct for the noise being absorbed in a single channel, as opposed to one of two channels. The final column corrects the noise to the expected noise on day 12, assuming the noise falls off with the same power law as measured here ($\kappa=0.63$). For CPDv1, we assume the data was taken on day 7.
@@
Detector Name Dimensions Holding Mass (g) Surface Area (cm2) Sidewall Area (cm2)
This work, 4 mm 1 cm2 $\times$ 4 mm HangingAnthony-Petersen et al. (2024) 0.932 3.6 1.6
This work, 1 mm 1 cm2 $\times$ 1 mm HangingAnthony-Petersen et al. (2024) 0.233 2.4 0.4
SPICE 1% Anthony-Petersen et al. (2025) 1 cm2 $\times$ 1 mm HangingAnthony-Petersen et al. (2024) 0.233 2.4 0.4
CPD v1 Fink et al. (2021); Alkhatib et al. (2021) 7.6 cm Ø$\times$ 1 mm Cirlex clamps 10.6 47.75 2.39
@@
CPD v2444Not previously published, an updated version of the detector described in Ref. Fink et al. (2021). 7.6 cm Ø$\times$ 1 mm Sapphire clamps 10.6 47.75 2.39
CRESST 0.35g Si 2022Angloher et al. (2023) (2 cm)2 $\times$ 0.4 mm Copper clamps 0.35 8.32 0.32
Table 3: Dimensions, weight, and holding of low threshold detectors considered in this work

### v3

@@ VI Acknowledgments
Supplementary Materials Section H Background Events and Correlated Noise Scaling With Mass vs. Surface Area, Comparison to Other Experiments
Using the 1 mm and 4 mm detectors characterized in this work, we show that both the rate of above threshold background events (i.e. LEE) and the correlated noise amplitude scale with the thickness of the detector substrate. This observation is compatible with three possible origins of such noise and backgrounds. First, these events and this noise could originate from the sidewalls of the detector, caused by e.g. saw damage from dicing. Alternatively, they could originate from the bulk volume of the detector substrate itself. Finally, the LEE rate could scale with mass of the substrate, but not originate in the volume. To our knowledge the only hypothesis that fits this third category is creep/plastic deformation of the structural wirebond foot that scales weight of the substrate.
The first two hypotheses could be distinguished by detectors with different sidewall area to volume ratios, testing whether the noise and backgrounds scale linearly with either detector volume or sidewall area. Unfortunately, no detectors with different sidewall area to volume ratios were manufactured at the same time as the detectors used in this work, preventing us from performing a direct test with the expectation of well controlled systematic variables. We aim to fabricate such detectors and directly test sidewall area vs. detector volume scaling of noise and background event rates in future work.
However, previously run detectors (the “CPDv1” detector in Fink et al. (2021) and an updated, previously unpublished version of the same detector, “CPDv2”) with different sidewall area to volume ratios are available for comparison. We compare these two form factors of detectors (1 cm2 for the 1 mm, 4 mm, and SPICE 1% detectors vs. the 76.2mm diameter by 1mm thick CPD detectors) in Fig. 16, and see that volume (as opposed to sidewall surface area) scaling appears to better match the data at higher energies. As the CPD detectors are single channel, higher threshold detectors, we expect that singles and noise triggers would contribute a significant number of events just above threshold, possibly explaining the difference in observed rate between the cm2 detectors and CPD detectors near threshold.
Figure 16: The rate of events observed in the 1 mm and 4 mm detectors compared to the SPICE 1% detectorAnthony-Petersen et al. (2025); Chang et al. (2025), the CPDv1 detectorAlkhatib et al. (2021); Fink et al. (2021), and the CPDv2 detector (previously unpublished) using the same coloring and energy range as in Fig. 5, scaling both by the mass of the detector (top) and by the sidewall area (bottom). As we do not know how the CRESST 0.35g Si detectorAngloher et al. (2023) was diced (i.e. which faces are sawed sidewalls vs. polished top and bottom faces), we do not include that detector for comparison. Note that the CPDv1 and v2 detectors are single channel detectors that cannot discriminate “singles,” possibly contributing to the rise in events just above threshold of those detectors. At higher energies, the LEE rate appears to better match between the two detector form factors (1 cm2 detectors: 1 mm, 4 mm, SPICE 1% detectors, vs. 3 inch diameter CPD detectors) when scaling by mass as opposed to sidewall area, suggesting that the LEE causing these high energy events originates within the bulk of the detector. A number of possible systematics effects might obscure an underlying scaling, see text for further discussion.
Unfortunately, several potential systematic effects make direct comparison with the 1 mm and 4 mm detectors worse than a better controlled test.
  • Time since cooldown: while in this work, we do not seem to observe a change in the observed LEE rate as a function of time above 5 eV, other detectors (see e.g. Ref. Angloher et al. (2023)) have observed a clear rate vs. time dependence at 10s of eV. Because the data for CPDv1 was taken significantly later in the run than the data for the CPDv2 and cm2 detectors (34 days vs. 11-12 days into the run, see table 2), rate vs. time dependencies could complicate the comparison of CPDv1 to the other detectors.
  • Detector holding: The 1 mm and 4 mm detectors as well as the SPICE 1% detector are suspended by wire bonds as described in Ref. Anthony-Petersen et al. (2024), while the CPDv1 and v2 detectors are clamped with Cirlex and sapphire clamps respectively. As detector holding methods which induce high stress in the detector are known to create backgrounds in the detector Astrom et al. (2006); Anthony-Petersen et al. (2024), the difference in holding methods between the two form factors of detectors could possibly add an additional LEE source in one or both sets of detectors. However, recent research by the CRESST collaboration led them to conclude that the LEE rate did not strongly vary between detectors with different clamping schemesFrancesca Pucci where the clamping was sufficiently weak such that it did not observavbly crack the detector substrate (as it did in Ref. Astrom et al. (2006)). This result is compatible with our observations assuming that the LEE rate scales with detector volume.
  • Time since manufacture: The time elapsed between detector fabrication and an LEE measurement could effect the measured LEE rate, for example if heating the detector substrate during e.g. photoresist baking changed the LEE rate. The estimated time between fabrication and data taking for each detector is summarized in Table 2. We see no clear relationship between the observed LEE rate and the time since fabrication, but leave open the possibility that such a systematic effect could obscure underlying trends in the data.
  • Single channel vs. two channel detectors: The CPD detectors are single channel detectors, and consequently cannot discriminate singles and shared events. The spectra for the CPD detectors therefore include singles, which presumably contribute significantly to the spectrum close to threshold, where the CPD detector spectra diverge from the cm2 detector spectra.
  • Differences in threshold and triggering: The CPD detectors have significantly higher thresholds than the cm2 detectors, presumably due in part to their phonon noise scaling with detector volume (as we observe in this work).
@@
Detector Name Dimensions Holding Mass (g) Surface Area (cm2) Sidewall Area (cm2) Day in Run Days Since Fab
This work, 4 mm 1 cm2 $\times$ 4 mm HangingAnthony-Petersen et al. (2024) 0.932 3.6 1.6 12 $\sim$ 90
This work, 1 mm 1 cm2 $\times$ 1 mm HangingAnthony-Petersen et al. (2024) 0.233 2.4 0.4 12 $\sim$ 90
SPICE 1% Anthony-Petersen et al. (2025) 1 cm2 $\times$ 1 mm HangingAnthony-Petersen et al. (2024) 0.233 2.4 0.4 11 $\sim$ 191
CPD v1 Fink et al. (2021); Alkhatib et al. (2021) 7.6 cm $\times$ 1 mm Cirlex clamps 10.6 47.75 2.39 34 $\sim$ 685
@@
CPD v2111Not previously published, an updated version of the detector described in Ref. Fink et al. (2021). 7.6 cm $\times$ 1 mm Sapphire clamps 10.6 47.75 2.39 12 $\sim$ 355
CRESST 0.35g Si 2022Angloher et al. (2023) (2 cm)2 $\times$ 0.4 mm Copper clamps 0.35 8.32 0.32 $\sim$250-350 Unknown
Table 2: Dimensions, weight, and holding of low threshold detectors considered in this work. “Day of Run” indicates how long after cooldown the spectra in Figs. 5, 16 were taken. “Days since Fab” estimates the elapsed time between the end of device fabrication and the start of datataking.
While we believe that it is relatively improbable that any of these systematics should dominate our backgrounds, and similarly improbable that such systematics would happen to coincidentally make backgrounds in the cm2 and CPD detectors appear to be compatible with volume scaling, we cannot eliminate the possibility that our observations are due to poorly controlled systematic effects. We aim to perform a more carefully controlled test of mass vs. sidewall area scaling with a future set of detectors.
The fact that the shared above threshold LEE seems consistent with mass/volume scaling for different structural support techniques (wirebond hanging, Cirlex clamping, sapphire clamping, copper clamping) disfavors the hypothesis that the above threshold shared LEE or phonon noise are due to wirebond foot creep.
Detector Name Correlated Noise Day of Run Noise / Mass Scaled Noise / Mass Noise / Sidewall Area Scaled Noise / Sidewall Area
(W2/Hz) (W2/(Hz g)) (W2/(Hz g)) (W2/(Hz cm2)) (W2/(Hz cm2))
This work, 4 mm $3.6\times 10^{-36}$ 12 $3.8\times 10^{-36}$ $3.8\times 10^{-36}$ $2.3\times 10^{-36}$ $2.3\times 10^{-36}$
This work, 1 mm $7.6\times 10^{-37}$ 12 $3.2\times 10^{-36}$ $3.2\times 10^{-36}$ $1.9\times 10^{-36}$ $1.9\times 10^{-36}$
SPICE 1%  [^2] This detector had degraded phonon collection efficiency and therefore lower observed noise due to accidental etching of its tungsten TESs. We correct for this by scaling the noise up by the ratio of the phonon collection efficiencies measured for this work and for the SPICE 1% detector, i.e. (39.1%/28.9%) = 1.33. Anthony-Petersen et al. (2025) $1.0\times 10^{-36}$ 7 $4.4\times 10^{-36}$ $3.1\times 10^{-36}$ $2.5\times 10^{-36}$ $1.8\times 10^{-36}$
CPD v1333Single channel detector, noise reduced by a factor of two for comparison to two channel detectors. Fink et al. (2021); Alkhatib et al. (2021) $4.2\times 10^{-35}$ 24 $4.0\times 10^{-36}$ $6.2\times 10^{-36}$ $1.7\times 10^{-35}$ $2.6\times 10^{-35}$
CPD v2444Not previously published, an updated version of the detector described in Ref. Fink et al. (2021). Single channel detector, noise reduced by a factor of two for comparison to two channel detectors. 4.7 $\times 10^{-35}$ 6 $4.5\times 10^{-36}$ $2.9\times 10^{-36}$ $2.0\times 10^{-35}$ $1.3\times 10^{-35}$
Table 3: Correlated noise amplitude comparison in different low threshold phonon detectors. The two single channel CPD detectors have their noise reduced by a factor of two to correct for the noise being absorbed in a single channel, as opposed to one of two channels. The final column scales the noise to the expected noise on day 12, assuming the noise falls off with the same power law as measured here ($\kappa=0.63$).